40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 29 14:43:12 2011, Koji, Update, LSC, LSC model updated 
    Reply  Wed Jul 6 23:03:57 2011, kiwamu, Update, LSC, modified locking scripts 
Message ID: 4912     Entry time: Wed Jun 29 14:43:12 2011     Reply to this: 4949
Author: Koji 
Type: Update 
Category: LSC 
Subject: LSC model updated 

LSC model has been updated and running,

- Now the power and signal recycling cavity lengths are named "PRCL" and "SRCL" in stead of three letter names without "L".

- Names for the trigger monitor were fixed. They are now "C1:LSC-DARM_TRIG_MON", etc., instead of "...NORM"

- Channel order of the DC signals for PDDC_MTRX and TRIG_MTRX were changed.

It was "TRX, TRY, REFL, AS, POP, POX, POY" but now "AS, REFL, POP, POX, POY, TRX, TRY".

We should change the locking script to accomodate these changes.

ELOG V3.1.3-